OpenStudy (anonymous):

the volume of a cylinder is 628cm^3. find the radius of the base if the cylinder has a height of 8cm.

OpenStudy (31356):

3. Here you're working backwards. The volume of a cylinder is (pi)*(radius^2)* height. So you have 628 = (pi)(r^2)(8) 25=r^2 5 cm = r

OpenStudy (31356):

Medal?

OpenStudy (anonymous):

im sorry, i'm kind of knew. how do i give you a medal?

OpenStudy (anonymous):

*new

OpenStudy (31356):

Thanks!

OpenStudy (anonymous):

your welcome!